Suppose that the nominal interest rate is 4.2%, the real interest rate is 2.8%, real GDP grows at 1%, and this year's money supply is $11.438B. To the nearest million, the size of next year's money supply will be $________B.
Answer #1

As from fisher effect

i= r + π

π = inflation rate = 4.2-2.8 = 1.4%

So from growth rate version of Quantity Theory of money

%∆ in M + %∆ in V = %∆ in Y + π

%∆ in M =1+ 1.4 = 2.4%

assuming V is Constant

so M in next year = 11.438 + 11.438*.024

= $ 11.712512 B